OpenStudy (anonymous):

Jesse worked 3 1/2 hours to build a brick wall. He built 2 2/3 feet of the wall each hour. How long is the wall he built? A. 9 1/3 feet B. 9 feet C. 7 feet D. 6 1/6 feet

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

you would multiply the number hours (3 1/2) by the rate (2 2/3)

jimthompson5910 (jim_thompson5910):

what do you get when you do this?

jimthompson5910 (jim_thompson5910):

keep in mind that 3 1/2 = 3 + 1/2 3 1/2 = 6/2 + 1/2 3 1/2 = 7/2

jimthompson5910 (jim_thompson5910):

and 2 2/3 = 2+2/3 2 2/3 = 6/3+2/3 2 2/3 = 8/3

jimthompson5910 (jim_thompson5910):

so (3 1/2) x (2 2/3) is the same as (7/2) x (8/3)

OpenStudy (anonymous):

9 1/3

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

yaay im done with math :) r you good at science? 6th grade science? lol sorry for bothering you :p

jimthompson5910 (jim_thompson5910):

not really, I'm ok at it so it's not as guaranteed with math

OpenStudy (anonymous):

can you try ? :)

jimthompson5910 (jim_thompson5910):

sure let's see what you got

OpenStudy (anonymous):

Directions: Read each step. Then put the steps in order from first to last. Write 1 for the first step, 2 for the second step, and so on. _____1. Fossil Fuels are burned. The thermal energy of the burning fuel turns water into steam. _____2. Over millions of years, the chemical energy in ancient organisms is transformed into the chemical energy of fossil fuels. _____3. The turbine turns a generator. The kinetic energy of the generator is converted to electrical energy. _____4. Organisms transform the radiant energy in sunlight into chemical energy. _____5. The kinetic energy of steam is transferred to a turbine

jimthompson5910 (jim_thompson5910):

ok one sec

OpenStudy (anonymous):

thanks :)

jimthompson5910 (jim_thompson5910):

ok this sounds like a time line

jimthompson5910 (jim_thompson5910):

from where energy originally comes from to where it ends up

jimthompson5910 (jim_thompson5910):

so you would start with the sun

jimthompson5910 (jim_thompson5910):

that energy gets into plants

jimthompson5910 (jim_thompson5910):

so that's why step 4 is the first step

OpenStudy (anonymous):

ok :)

jimthompson5910 (jim_thompson5910):

i mean, this is the first step _____4. Organisms transform the radiant energy in sunlight into chemical energy.

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

energy comes from the sun ---> goes to plants then the plants die and turn into fossils

jimthompson5910 (jim_thompson5910):

even though when you think of fossils as dinosaurs, the majority of fossils are plants and tiny creatures (since there were a lot more plants than dinosaurs)

jimthompson5910 (jim_thompson5910):

so that's why _____2. Over millions of years, the chemical energy in ancient organisms is transformed into the chemical energy of fossil fuels. is next

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

number 1 would be step 3?

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

_____1. Fossil Fuels are burned. The thermal energy of the burning fuel turns water into steam. is next

jimthompson5910 (jim_thompson5910):

then what?

OpenStudy (anonymous):

number 3 would be step 4?

OpenStudy (anonymous):

NOO

OpenStudy (anonymous):

number 5 would be step 4

jimthompson5910 (jim_thompson5910):

yep, 5, then 3

OpenStudy (anonymous):

low charge on my computer :S

jimthompson5910 (jim_thompson5910):

speaking of electrical energy lol

OpenStudy (anonymous):

lol!! ok thank you soo much

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

byee :) btw my sister @gabylovesyou recommended me with u

jimthompson5910 (jim_thompson5910):

ah that name sounds familiar lol

jimthompson5910 (jim_thompson5910):

have a good night

OpenStudy (anonymous):

lol!!! :) nighty night night

jimthompson5910 (jim_thompson5910):

good night